OpenStudy (df001):

Find the exact value of sinΘ, cosΘ, and tanΘ using reference angles. Θ=330 degrees

OpenStudy (df001):

Help how do i do this!??! I need quick help please!

OpenStudy (mayankdevnani):

\[\large \bf \sin330^o=\sin(360-30)=-\sin(30^o)=-\frac{1}{2}\]

OpenStudy (mayankdevnani):

similarly, you can find for cos and tan

OpenStudy (lochana):

 this is helpful for your problem

OpenStudy (triciaal):